a motorboat can maintain a constant speed of 16 miles per hour relative to the water. The boat makes a trip upstream to a certain point in 20 minutes the return takes 15 minutes. what is the speed of the current?
Answer by stanbon(75887)

a motorboat can maintain a constant speed of 16 miles per hour relative to the water. The boat makes a trip upstream to a certain point in 20 minutes the return takes 15 minutes. what is the speed of the current?
-----------------------
Upstream DATA:
time = 1/3 hr ; rate = 16-c mph ; distance = r*t = (1/3)(16-c) miles
-----
Downstream DATA:
time = 1/4 hr ; rate = 16+c mph ; distance = r*t = (1/4)(16+c)
--------
Equation:
distance = distance
(1/3)(16-c) = (1/4(16+c)
----
4(16-c) = 3(16+c)
64-c = 48+c
16 = 2c
current = 8 mph
=================
Cheers,
Stan H.
=================